OpenStudy (anonymous):

If the midpoint between (x, 3) and (9, 9) is (7, 6), find the value of x. Answer -2 23 5 9

OpenStudy (jdoe0001):

\(\bf \text{middle point of 2 points }\\ \left(\cfrac{x_2 + x_1}{2}\quad ,\quad \cfrac{y_2 + y_1}{2} \right)\\ \quad \\ (x,3)\qquad (9,9)\\ \quad \\ \left(\color{blue}{\cfrac{9 + x}{2}}\quad ,\quad \cfrac{9 + 3}{2} \right)\implies (\color{blue}{7},6)\qquad thus\\ \quad \\ \cfrac{9 + x}{2}=7 \) solve for "x"

OpenStudy (anonymous):

ok 9+x=14 x=5?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

thanks <3 can you help me with 2 more?

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

M is the midpoint of . The coordinates of A are (2,3) and the coordinates of M are (4.5,6). Find the coordinates of B. Answer (3.25, 4.5) (7, 9) (2.5, 3) (11, 15)

OpenStudy (anonymous):

M is the midpoint of segment AB*

OpenStudy (anonymous):

actually I tried it, is it 3.25,4.5?

OpenStudy (jdoe0001):

hmm one sec

OpenStudy (jdoe0001):

I was a bit caught... up lemme set it

OpenStudy (jdoe0001):

\(\bf (2,3)\qquad (x_2 ,y_2 )\\ \quad \\ \left(\cfrac{x_2 + 2}{2}\quad ,\quad \cfrac{y_2 + 3}{2} \right)\implies (4.5,6)\\ \quad \\ \cfrac{x_2 + 2}{2}=4.5\qquad \qquad \cfrac{y_2 + 3}{2}=6\)